# Speech-FT: Merging Pre-trained And Fine-Tuned Speech Representation Models For Cross-Task Generalization

### v2

@@ VI-A Model Merging
Model merging is a technique that combines multiple models into a single model while preserving the key attributes of each original model [16]. Weight-space interpolation can be viewed as a form of model merging, where a pre-trained and a fine-tuned model are combined in parameter space.
Numerous studies have explored model merging in computer vision, including techniques for multi-target domain adaptation [49], efficient integration of vision-language models using reusable modules [50], and maintaining task-specific information through selective parameter merging [51]. In natural language processing, model merging methods have been applied to mitigate the alignment tax of RLHF via model averaging [52], enhance multi-task learning with large language models [53], and adapt models to low-resource languages by combining source and target models [54].
In the speech domain, several studies have investigated the applications of model merging. These include merging ASR models trained on different dataset distributions to enhance out-of-distribution generalization [55, 56], merging TTS models to interpolate speaker attributes [57], merging speech translation models to expand language pairs [58], and merging fine-tuned ASR models to support lifelong learning of unseen accents [22].
@@ VI-B Out-of-Distribution Generalization
Out-of-distribution (OOD) generalization aims to preserve a model’s performance on OOD datasets after fine-tuning. Several studies in computer vision have explored OOD generalization for vision pre-trained models that could perform zero-shot inference. For instance, LP-FT [11] introduces a simple two-step approach—linear probing followed by full fine-tuning—to mitigate distortion in pre-trained features by first learning a near-optimal linear head during the initial steps. Model Soup [59] demonstrates that averaging the weights of models fine-tuned with different hyperparameter configurations can improve robustness on OOD datasets. Wise-FT [19] proposes merging the weights of pre-trained and fine-tuned models to enhance robustness on OOD datasets.
A closely related approach to Speech-FT is Wise-FT [19], which also employs weight-space interpolation between pre-trained and fine-tuned models. However, it differs in several key aspects. First, Wise-FT is specifically designed to enhance out-of-distribution (OOD) generalization in image classification, where the model is tested under distributional shifts such as changes in style or image perturbations. In contrast, Speech-FT addresses a more complex challenge—preserving generalization across diverse downstream tasks in speech representation learning. Second, Speech-FT introduces a Stable-FT stage before interpolation, which is shown in Section V-B to be critical for maintaining cross-task generalization. Notably, using weight-space interpolation alone may be insufficient to improve speech representation. Third, we evaluate the effectiveness of Speech-FT across diverse fine-tuning tasks in speech processing. We also evaluate Speech-FT under unsupervised language adaptation and in scenarios involving multiple fine-tuning tasks. These explorations have not been addressed in prior works. Finally, our study provides an analysis of why weight-space interpolation is effective in speech representation learning, offering insights that are absent from prior works.
@@ VI-C Continual Learning
Continual learning, or lifelong learning, enables models to adapt to new domain data while retaining previously acquired knowledge. It addresses the issue of catastrophic forgetting [12] and is essential for handling dynamic and evolving environments. Continual learning typically involves multiple rounds of fine-tuning to simulate real-world scenarios where models continually learn from new data.
Several prior studies have explored continual learning for automatic speech recognition [22, 60, 61]. While we also aim to preserve previously acquired information, our objective is fundamentally different from that of continual learning. Rather than simulating a long sequence of tasks, we focus primarily on enhancing speech representations after a single fine-tuning stage. Since our main goal is to improve representation quality—rather than to ensure robustness under sequential task adaptation—it is unnecessary to introduce task sequences into our setting.
@@ VI-D Fine-tuning Speech Representation Models
Several prior works have attempted to fine-tune speech representation models, with most focusing on improving performance on specific downstream tasks, such as speech recognition [62, 63, 64, 48] and emotion recognition [65, 66]. In contrast, our work falls within the domain of speech representation learning—we aim to improve the overall quality of speech representations through fine-tuning, with a focus on enhancing performance across a broad range of speech-related tasks.
One prior study explored improving speech representations via supervised fine-tuning tasks [9], but it did not explicitly address the preservation of cross-task generalization. As a result, their approach (denoted as regular fine-tuning in Table I) fails to generalize well across tasks.
Another study investigated continuing the pre-training process on a low-resource language [10], but their focus was on improving performance in that specific language. In contrast, our goal is to preserve generalization across both previously learned and new languages.

### v3

@@ VI-A Model Merging
Model merging is a technique that combines multiple models into a single model while preserving the key attributes of each original model [ilharco2022editing]. Weight-space interpolation can be viewed as a form of model merging, where a pre-trained and a fine-tuned model are combined in parameter space.
Numerous studies have explored model merging in computer vision, including techniques for multi-target domain adaptation [li2024training] and maintaining task-specific information through selective parameter merging [marczak2024magmax]. In natural language processing, model merging methods have been applied to mitigate the alignment tax of RLHF via model averaging [lin2024mitigating] and enhance multi-task learning with large language models [zhou2024metagpt].
In the speech domain, several studies have investigated the applications of model merging. These include merging ASR models trained on different dataset distributions to enhance out-of-distribution generalization [ramesh2024task, plantinga2024parameter], merging TTS models to interpolate speaker attributes [murata2024attribute], merging speech translation models to expand language pairs [cheng2024task], and merging fine-tuned ASR models to support lifelong learning of unseen accents [kulshreshtha2024sequential].
@@ VI-B Out-of-Distribution Generalization
Out-of-distribution (OOD) generalization aims to preserve a model’s performance on OOD datasets after fine-tuning. Several studies in computer vision have explored OOD generalization for vision pre-trained models that could perform zero-shot inference. For instance, LP-FT [kumar2022fine] introduces a simple two-step approach—linear probing followed by full fine-tuning—to mitigate distortion in pre-trained features by first learning a near-optimal linear head during the initial steps. Model Soup [wortsman2022model] demonstrates that averaging the weights of models fine-tuned with different hyperparameter configurations can improve robustness on OOD datasets.
A relevant approach to Speech-FT in a different context is Wise-FT [wortsman2022robust], which also employs weight-space interpolation between pre-trained and fine-tuned models. However, it differs in several key aspects. First, Wise-FT is specifically designed to enhance out-of-distribution (OOD) generalization in image classification, where the model is tested under distributional shifts such as changes in style or image perturbations. In contrast, Speech-FT addresses a more complex challenge—preserving generalization across diverse downstream tasks in speech representation learning. Second, Speech-FT introduces a Stable-FT stage before interpolation, which is shown in Section V-B to be critical for maintaining cross-task generalization. Notably, using weight-space interpolation alone may be insufficient to improve speech representation. Third, we evaluate the effectiveness of Speech-FT across diverse fine-tuning tasks in speech processing. We also evaluate Speech-FT under unsupervised language adaptation and in scenarios involving multiple fine-tuning tasks. These explorations have not been addressed in prior works. Finally, our study provides an analysis of why weight-space interpolation is effective in speech representation learning, offering insights that are absent from prior works.
@@ VI-C Continual Learning
Continual learning, or lifelong learning, enables models to adapt to new domain data while retaining previously acquired knowledge. It addresses the issue of catastrophic forgetting [kirkpatrick2017overcoming] and is essential for handling dynamic and evolving environments. Continual learning typically involves multiple rounds of fine-tuning to simulate real-world scenarios where models continually learn from new data.
Several prior studies have explored continual learning for automatic speech recognition [kulshreshtha2024sequential, houston2020continual]. While we also aim to preserve previously acquired information, our objective is fundamentally different from that of continual learning. Rather than simulating a long sequence of tasks, we focus primarily on enhancing speech representations after a single fine-tuning stage. Since our main goal is to improve representation quality—rather than to ensure robustness under sequential task adaptation—it is unnecessary to introduce task sequences into our setting.
@@ VI-D Fine-tuning Speech Representation Models
Several prior works have attempted to fine-tune speech representation models, with most focusing on improving performance on specific downstream tasks, such as speech recognition [yang2024finetuning, chen2023explore], emotion recognition [chen2023emotionfinetune], and anti-spoofing detection [pan24c_interspeech, guragain2024speech]. In contrast, our work falls within the domain of speech representation learning—we aim to improve the overall quality of speech representations through fine-tuning, with a focus on enhancing performance across a broad range of speech-related tasks.
One prior study explored improving speech representations via supervised fine-tuning tasks [chen2021speech], but it did not explicitly address the preservation of cross-task generalization. As a result, their approach fails to generalize well across tasks.
Another study investigated continuing the pre-training process on a low-resource language [getman2024happens], but their focus was on improving performance in that specific language. In contrast, our goal is to preserve generalization across both previously learned and new languages.